OpenStudy (anonymous):

What value of k makes the statement true? –(k + 12) = 6 + (–12) A. –18 B. –6 C. 6 D. 18

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

@vampiregodess

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@angel12310

OpenStudy (anonymous):

i say a

OpenStudy (jennyrlz):

hello ^.^

OpenStudy (jennyrlz):

it isnt A

OpenStudy (anonymous):

aww hold on

OpenStudy (mathmate):

If you think its A, and would like to check the answer, the best way is to put the value into the equation and evaluate to see if both sides are equal!

OpenStudy (jennyrlz):

–(k + 12) = 6 + (–12)

OpenStudy (jennyrlz):

lets work this out

OpenStudy (anonymous):

ok

OpenStudy (jennyrlz):

what is 6 -12?

OpenStudy (jennyrlz):

PS: if it where 6-(-12) the double negatives would turn into a possitive

OpenStudy (anonymous):

6

OpenStudy (jennyrlz):

-6*

OpenStudy (anonymous):

+

OpenStudy (anonymous):

thx

OpenStudy (jennyrlz):

np

OpenStudy (jennyrlz):

so now we have -(k+12) = -6

OpenStudy (jennyrlz):

now we get rid of those pesky negatives

OpenStudy (jennyrlz):

instead of distributing that negative i am just gonna take it out of both sides

OpenStudy (jennyrlz):

so now we have k + 12 = 6

OpenStudy (anonymous):

THX U SO MUCH @jennyrlz

OpenStudy (jennyrlz):

:) np

OpenStudy (jennyrlz):

did you get the answer?

OpenStudy (jennyrlz):

yea :), btw you couldve also distruibuted the negative i just found it easier this way

OpenStudy (anonymous):

oh ok yayy so b right

OpenStudy (jennyrlz):

for example -(k+12)=-6 -k -12 = - 6 -k= 6 get rid of the negative by multiplying both sides by the negative one ( switchign the signs same way as the first) k = -6 B

OpenStudy (jennyrlz):

and yup B :)

OpenStudy (anonymous):

wow thx so much and 4 all ur help and tips thxxxxxxx

OpenStudy (jennyrlz):

np :P

OpenStudy (jennyrlz):

glad i could help xD